OpenStudy (anonymous):

what is the units digit of 7^282

OpenStudy (mathstudent55):

7^2 ends in 9 7^12 ends in 1 7^22 ends in 9 7^32 ends in 1 7^42 ends in 9 7^52 ends in 1 etc.

OpenStudy (anonymous):

Ans is 9 Explanation U see 7 has a cycle of 4 for units digit For Ex: 7^1=7 7^2=49 (9 is units digit) 7^3=343 (3 being units digit) 7^4=2401(1 being units digit) on 7^5 last digit will be 7 again so u first divide the exponent 7^Y by 4 (Here y=282) Remainder will be left power of 7 (Here 7^2=49) Hence 9 will be the units digit

OpenStudy (anonymous):

Thank you so much! You were helpful. Many people just type the answer and don't explain... You did